The hydronium ion concentration in an aqueous solution at 25°C is 3.1×10-2M.

The hydroxide ion concentration is  M.

The pH of this solution is .

The pOH is
Answer #1

hydronium ion concentration

[H3O+] = 3.1 x 10^-2 M

pH = - log[H3O+]

= - log (3.1*10^-2)

= 1.51

pOH = 14 - 1.51 = 12.49

[OH-] = 10^-pOH = 10^-12.49 = 3.24 x 10^-13 M